OpenStudy (anonymous):

Which formula would you use when you are looking for the slope of a line with two points

OpenStudy (texaschic101):

you would use the slope formula : slope = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

Yep. You use the slope formula. For example: 3,-6 6,12 y2-y1=12-(-6)=18 x2-x1=6-3=3 slope or m=18/3=6/1=6